QD=100 000-5p




QD=-25000+20p




What is the price




What is the quantity demanded




What is the quantity supplies

1
Expert's answer
2022-05-09T14:51:38-0400

"Q_d=100000-5p,"

"Q_s=-25000+20p,"

"Q_d=Q_s,"

"100000-5p=-25000+20p,"

"p=5000,"

"Q_d=Q_s=75000."